# RLSA-PFL: Robust Lightweight Secure Aggregation with Model Inconsistency Detection in Privacy-Preserving Federated Learning

### v2

@@ 2.1 DP, HE, and SMC Based Schemes
In DP-based secure aggregation schemes, random noise is added to users’ gradients before they are sent to the aggregator server [22]. This process protects the sensitive information within the gradients by making it difficult to reverse-engineer the original values, as DP introduces controlled noise to the data. Despite this added layer of privacy, the server can still aggregate these perturbed gradients to approximate the true model updates, owing to the mathematical properties of DP. However, a key challenge with DP is the inevitable trade-off between the level of privacy and the utility of the data. The more noise that is introduced to protect privacy, the less accurate the aggregated model becomes, which can potentially lead to significant degradation in model performance [17]. Striking a delicate balance between safeguarding individual privacy and maintaining data utility is a central concern in the application of DP in FL, and finding the optimal point on this spectrum remains an area of active research. Some notable works in this domain include [22, 23, 24].
@@
In Homomorphic Encryption (HE)-based aggregation schemes, such as those proposed in [25, 26, 27, 28, 29] users first apply computation-intensive algorithms to encrypt their local model parameters before sending them to the central server. HE allows for arithmetic operations to be performed directly on encrypted data without needing to decrypt it first. This means that the server can aggregate the encrypted local model parameters by performing operations like addition or multiplication directly on the ciphertext. After the aggregation process, the server either sends the aggregated encrypted result back to the users for decryption or continues the training process directly on the ciphertext. This approach ensures that sensitive gradient information remains protected throughout the entire process, as the data remains encrypted during both transmission and computation. However, the use of HE introduces significant computational overhead, both in terms of the initial encryption performed by the participants and the subsequent operations carried out by the server on the encrypted data [17], [1].
@@
SMC is another cryptographic technique that has been applied in secure aggregation within FL [30, 21, 31]. SMC allows multiple participants, each with their own private data, to collaboratively compute a desired objective function without revealing their data to others. This method ensures that each participant’s data remains confidential while still enabling the computation of an accurate result [11]. However, SMC-based schemes tend to be inefficient, as they often involve significant communication overhead and face challenges in managing user dropouts [19], [1].
2.2 Masking-Based Schemes
@@
The core idea of masking-based schemes is to secure users’ local model parameters by adding random values (often referred to as one-time pads) to them before they are sent to the aggregator server [12]. These random values are designed so that they cancel out during the aggregation process, allowing the aggregator server to recover the aggregated plaintext gradient values for all participants in that round. The primary goal of the masking-based approach is to protect individual users’ local model parameters from being exposed to unauthorized parties while still enabling the aggregator server to access the aggregated gradient in its plaintext form [1]. Our work aligns with this approach, and we provide a detailed literature review on it below.
@@
In [12], Bonawitz et al. proposed a practical secure aggregation scheme using masking to conceal individual users’ local model parameters from the aggregator server. The scheme also employs secret sharing techniques to accommodate user dropouts, ensuring that the learning process is not affected by these dropouts. However, while it works well for one round of training, it becomes highly inefficient when multiple training rounds are required (which is essential for most real-world FL applications) due to the need for an expensive setup phase that involves four communication rounds to establish shared randomness and pairwise keys in every FL training round. Subsequently, several works have been conducted to improve the security and efficiency of [12] such as [18, 16, 15, 17, 13, 14]. In [18] and [16], the authors proposed two schemes to add verifiability on top of the protocol [12]. In [15], Fereidooni et al. used secret sharing and homomorphic encryption to achieve secure aggregation without relying on a trusted third party to generate any public/private key pair for the clients. In [17], Liu et al. applied the concepts of homomorphic pseudo-random generator and Shamir Secret Sharing technique in achieving user dropouts and reducing the communication costs among the users, which is secure against both semi-honest and malicious adversaries. In [13], the authors reduced the communication overhead of [12] by utilizing a logarithmic degree k-regular graph. In [32], Liu et al. proposed a Dynamic User Clustering scheme that builds on existing masking-based secure aggregation schemes, such as [12], and incorporates a sparsification technique to address the interoperability issues with sparsification. In [33], Fu et al. proposed a blockchain-based decentralized secure aggregation scheme to replace the central aggregator server. The scheme uses masking and Shamir’s Secret Sharing to ensure privacy. However, schemes such as [18, 16, 15, 17, 13, 32, 33] can only accommodate limited user dropouts due to their reliability in the secret sharing method. Further, the users cannot join in real-time, thereby limiting their flexibility in real-world applications. Moreover, each client must undergo pairwise random seed negotiation, share computation, and data transmission in each round of training, causing the system’s complexity to escalate significantly as the number of clients grows. In [14], Eltaras et al. proposed a pairwise masking-based secure aggregation scheme that uses auxiliary nodes to achieve verifiability and handle user dropouts. However, this scheme requires key agreement among users and auxiliary nodes, which increases communication overhead. Additionally, it does not address the malicious security model, where the server itself may behave maliciously. In [19], Ma et al. proposed “Flamingo”, a multi-round single server secure aggregation scheme that does not require an initialization setup phase for each training round (a single setup phase is sufficient). However, Flamingo cannot handle a dynamic environment where the users should be able to join the training session dynamically. Flamingo requires to know the participating users before the training starts. Moreover, Flamingo can tolerate up to one-third of corrupted users, which includes both regular and decryptor users, among the total users. In [34], Wang et al. proposed a single mask scheme based on the Decisional Composite Residuosity (DCR) assumption and used non-interactive zero-knowledge (NIZK) proofs to achieve result verification. However, it uses computationally expensive bilinear pairing operations. In [35], Yang et al. proposed a single mask secure aggregation scheme for FL by combining the concepts of homomorphic Pseudorandom Generator, homomorphic Shamir secret sharing, and Paillier encryption. However, it can only tolerate $50\%$ of the user dropouts. In [36], Khojir et al. introduced a secure aggregation scheme based on additive secret sharing. The scheme utilizes a three-layered architecture (i.e., clients, middle servers, and lead server), which reduces communication costs for users compared to other schemes that rely on masking processes using secret sharing, which requires distributing secret shares among users, such as those in [18, 16, 15, 17, 13, 14]. However, it is unable to handle user dropouts, which is essential for real-world applications.
Table I compares the computation and communication overhead, security, and functionality of our scheme against notable works in masking-based secure aggregation, including SecAgg [12], SecAgg+ [13], FastSecAgg [21], EDRAgg [17], and Flamingo [19]. Our scheme offers better performance in both communication and computation costs at the user and aggregator sides. While our approach introduces intermediate servers to achieve enhanced security and functionality, this addition does not negatively impact the system’s overall performance or accuracy. In fact, these intermediate servers reduce the computational and communication burden on users, which will be further elaborated in the upcoming sections. Regarding security, our scheme supports a malicious threat model, just like SecAgg [12], SecAgg+ [13], EDRAgg [17], and Flamingo [19], which is a stronger security assumption compared to the semi-honest model. Additionally, our scheme demonstrates strong resilience to user dropouts, tolerating up to $m-2$ user dropouts, which is higher than many existing schemes. Like Flamingo [19], our scheme only requires one setup phase for the entire training session, significantly reducing communication overhead for users. Furthermore, both our scheme and Flamingo are capable of detecting model inconsistency attacks through model parameter verification techniques. Another benefit is that our scheme involves only two communication rounds between the user and the aggregator per training round, further minimizing communication costs compared to other methods.
3 Cryptographic Primitives
@@
In this section, we briefly introduce the core concepts of symmetric homomorphic encryption and digital signature, which are integral to the design of our scheme.
3.1 Symmetric Homomorphic Encryption
@@
Our scheme employs a key negation method, similar to that in [37], to mask the user’s local model parameters. This method is based on the symmetric homomorphic encryption mechanism proposed by [38]. In this section, we will briefly describe the work of [38]. The key negation method will be explained in Section 5.1.
@@
[…]
$=$ $m_{i}+m_{j}.$ (3)
The scheme described above has additive homomorphic properties [38]. The addition of two ciphertexts is illustrated in Equation 3.1. This property can also be extended to aggregate a set of ciphertexts using the corresponding aggregated secret keys, as shown in Equation 4.
$\sum_{1\leq i\leq n}m_{i}=$ $\mathsf{Dec}_{\sum_{1\leq i\leq n}k_{i}}\left(\sum_{1\leq i\leq n}c_{i}\right)$ (4)
$=$ $\sum_{1\leq i\leq n}c_{i}-\sum_{1\leq i\leq n}k_{i}\mod{q}.$ (5)
The described symmetric homomorphic encryption scheme is semantically secure (IND-CPA) if the secret keys $\mathtt{k_{i}}$, where $i\in\{1,2,\cdots,n\}$, are generated randomly and no keys are reused. Please refer to [38] for detailed security proof.
3.2 Signature Scheme
Our scheme employs digital signatures to ensure the integrity and authenticity of messages exchanged between entities, leveraging a UF-CMA secure scheme as described in [12]. In general, a digital signature scheme consists of the following probabilistic polynomial-time (PPT) algorithms:
@@ 3.2 Signature Scheme
Figure 1: System Architecture
4 System Architecture, Threat Model, Assumptions, and Goals of Our Scheme
In this section, we present the system architecture, threat model, and security assumptions, as well as the functionality and security goals of our proposed scheme. We first start with the system architecture.
4.1 System Architecture
Figure 1 illustrates the system architecture of our scheme, which comprises three primary entities: users, intermediate servers, and the central server which we termed as Aggregator.
  • Users are the edge devices, such as smartphones and IoT devices, for cross-device FL [^2] Cross-device FL refers to a distributed machine learning method where many devices, such as smartphones or IoT devices, train a shared global model without sharing their local data., or organizations like hospitals, banks, universities, and government agencies for cross-silo FL [^3] Cross-silo FL involves a small number of organizations, such as hospitals or businesses, collaboratively training a model without sharing their private data. Unlike cross-device FL, cross-silo FL usually deals with higher-quality data from trusted institutions and is designed for scenarios with stricter privacy and security requirements.. These users generate and own data locally. They perform local model training on their own data, downloading the current global model from the aggregator, training it, and then sending the updated model parameters to the intermediate servers and aggregator. Before transmission, the parameters are masked, ensuring that only the aggregator can recover the final aggregated model parameter, thereby maintaining the confidentiality of individual users’ local model data.
  • Intermediate servers act as aggregation points between users and the main aggregator. They are to reduce communication overhead by combining model updates from multiple users before sending them to the aggregator. These servers receive model updates from users, aggregate the updates, and then forward the aggregated results to the aggregator for final processing. Additionally, they can assist users in detecting model inconsistency attacks originating from the aggregator. In practical applications, fog nodes, edge servers, and third-party cloud services can function as intermediate servers, especially in edge computing or distributed system scenarios. By facilitating localized processing, these components reduce latency and bandwidth usage.
  • Aggregator is the central server that handles the entire FL process. It maintains the global model, coordinates the FL training rounds, and aggregates model updates directly from the intermediate servers and users. The aggregator initiates the training process by distributing the initial global model to users. After receiving aggregated updates from the intermediate servers, the aggregator combines these updates to improve the global model. It then sends the updated global model back to users for the next round of training.
4.2 Threat Model
@@
In the malicious model, we assume a more challenging scenario where both the aggregator and the intermediate servers may act maliciously, attempting to compromise the individual users’ local training models. In this scenario, we assume that the malicious aggregator and the malicious intermediate servers can collude with up to $m-2$ users, where $m$ is the total number of participating users in an FL round. Similarly, we assume that the malicious aggregator can collude with up to $n-1$ malicious intermediate servers in an FL round, where $n$ is the total number of participating intermediate servers. Additionally, our threat model considers model inconsistency attacks by the malicious aggregator, where the malicious aggregator provides different parameters to different users to exploit behavioral differences in the model updates, thereby inferring information on users’ datasets [20]. Our aim is to detect such behavior by the malicious aggregator.
@@
We also consider the possibility of malicious users in our threat model. These users can collude with other users as well as with the aggregator and the intermediate servers to gain knowledge of other users’ local training models. However, we assume that up to $m-2$ users can collude, meaning at least two users must remain honest at all times. Without this assumption, the users could collude with the aggregator to obtain the local training model update of a targeted user.
4.3 Assumption
We have made several assumptions in designing our scheme, which are listed below:
  • The aggregator and intermediate servers remain online at all times.
  • The aggregator and intermediate servers are aware of the participating users before initiating the FL process.
  • Each entity possesses a private and public key through a Public Key Infrastructure (PKI).
  • All users are identified by a unique identifier.
  • Each participating user has knowledge of the online intermediate servers.
  • An authenticated and private channel exists between users and intermediate servers, users and the aggregator, and between intermediate servers and the aggregator.
4.4 Goals
In this section, we present some of the primary security and functionality goals of our scheme.
  • Local Model Parameter Privacy: The primary objective of our scheme is to safeguard the privacy and confidentiality of individual users’ local model parameters. This ensures that no entity can infer or deduce these local model parameters, except for the aggregated global model that is shared with the aggregator.
  • Flexible User Dropout: Our scheme supports dynamic user participation in the training process. Users can join or leave at any FL round due to issues like connectivity, device limitations, or power constraints. Despite this, the global model continues to train effectively. It remains robust and converges, even if some users fail to provide their local updates.
  • Model Inconsistency Attack Detection: Our scheme is designed to detect any malicious behavior by the aggregator attempting to exploit the model inconsistency attack, as demonstrated in [20]. Our scheme addresses this vulnerability by ensuring that any such malicious tampering by the aggregator is promptly identified, maintaining the integrity of the FL process.
  • Elimination of Inter-User Communication: Our scheme aims to eliminate any interaction among users, which helps reduce both communication and computational overhead, in contrast to the approach used by [12] and its variants.
Figure 2: Sample Key Negation Mechanism

### v3

@@ 2.1. DP, HE, and SMC Based Protocols
In DP-based secure aggregation protocols, random noise is added to users’ gradients before they are sent to the aggregator server (Wei et al., 2020). This process protects the sensitive information within the gradients by making it difficult to reverse-engineer the original values, as DP introduces controlled noise to the data. Despite this added layer of privacy, the server can still aggregate these perturbed gradients to approximate the true model updates, owing to the mathematical properties of DP. However, a key challenge with DP is the inevitable trade-off between the level of privacy and the utility of the data. The more noise that is introduced to protect privacy, the less accurate the aggregated model becomes, which can potentially lead to significant degradation in model performance (Liu et al., 2023). Striking a delicate balance between safeguarding individual privacy and maintaining data utility is a central concern in the application of DP in FL, and finding the optimal point on this spectrum remains an area of active research. Some notable works in this domain include (Wei et al., 2020; Wang et al., 2020; Zhou et al., 2022).
@@
In HE-based aggregation protocols, such as those proposed in (Chan et al., 2012; Phong et al., 2018; Ma et al., 2022; et al., 2021b; Zhang et al., 2020) users first apply computation-intensive algorithms to encrypt their local model parameters before sending them to the central server. HE allows for arithmetic operations to be performed directly on encrypted data without requiring it to be decrypted first. This means that the server can aggregate the encrypted local model parameters by performing operations like addition or multiplication directly on the ciphertext. After the aggregation process, the server either sends the aggregated encrypted result back to the users for decryption or continues the training process directly on the ciphertext. This approach ensures that sensitive gradient information remains protected throughout the entire process, as the data remains encrypted during both transmission and computation. However, the use of HE introduces significant computational overhead, both in terms of the initial encryption performed by the participants and the subsequent operations carried out by the server on the encrypted data (Liu et al., 2023), (Liu et al., 2022).
@@
SMC is another cryptographic technique that has been applied in secure aggregation within FL (So et al., 2021; Kadhe et al., 2020; So et al., 2022). SMC allows multiple participants, each with their private data, to collaboratively compute a desired objective function without revealing their data to others. This method ensures that each participant’s data remains confidential while still enabling the computation of an accurate result (Yin et al., 2021). However, SMC-based protocols tend to be inefficient, as they often involve significant communication overhead and face challenges in managing user dropouts (Liu et al., 2022), (Ma et al., 2023).
2.2. Masking-Based Protocols
@@
The core idea of masking-based protocols is to secure users’ local model parameters by adding random values (often referred to as one-time pads) to them before they are sent to the aggregator server (et al., 2017). These random values are designed so that they cancel out during the aggregation process, allowing the aggregator server to recover the aggregated plaintext gradient values for all participants in that round. The primary goal of the masking-based approach is to protect individual users’ local model parameters from being exposed to unauthorized parties while still enabling the aggregator server to access the aggregated gradient in its plaintext form (Liu et al., 2022). Our work aligns with this approach.
@@
In (et al., 2017), Bonawitz et al. proposed a practical secure aggregation protocol that uses masking to conceal individual users’ local model parameters from the aggregator server. The protocol also employs secret sharing techniques to accommodate user dropouts, ensuring that the learning process remains unaffected. Building on the core concepts introduced by Bonawitz et al., several subsequent works have been proposed over the years to enhance security and functionality, including (Xu et al., 2020; Guo et al., 2021; et al., 2021a; Liu et al., 2023; Bell et al., 2020; Eltaras et al., 2023). In (Xu et al., 2020) and (Guo et al., 2021), the authors proposed two protocols to add verifiability on top of the protocol (et al., 2017). In (et al., 2021a), Fereidooni et al. used secret sharing and homomorphic encryption to achieve secure aggregation without relying on a trusted third party to generate any public/private key pair for the clients. In (Liu et al., 2023), Liu et al. applied the concepts of homomorphic pseudo-random generator and Shamir Secret Sharing technique in achieving user dropouts and reducing the communication costs among the users, which is secure against both semi-honest and malicious adversaries. In (Bell et al., 2020), the authors reduced the communication overhead of (et al., 2017) by utilizing a logarithmic degree k-regular graph. In (et al., 2024), Liu et al. proposed a Dynamic User Clustering protocol that builds on existing masking-based secure aggregation protocols, such as (et al., 2017), and incorporates a sparsification technique to address the interoperability issues with sparsification. In (Fu et al., 2024), Fu et al. proposed a blockchain-based decentralized secure aggregation protocol to replace the central aggregator server. The protocol uses masking and Shamir’s Secret Sharing to ensure privacy. In (Eltaras et al., 2023), Eltaras et al. proposed a pairwise masking-based secure aggregation protocol that uses auxiliary nodes to achieve verifiability and handle user dropouts. In (Yang et al., 2023), Yang et al. proposed a single mask secure aggregation protocol for FL by combining the concepts of homomorphic Pseudorandom Generator, homomorphic Shamir secret sharing, and Paillier encryption. However, all these protocols ((et al., 2017; Xu et al., 2020; Guo et al., 2021; et al., 2021a; Liu et al., 2023; Bell et al., 2020; Yang et al., 2023; Eltaras et al., 2023; et al., 2024; Fu et al., 2024)) work well for one round of training, they become highly inefficient when multiple training rounds are required (which is essential for most real-world FL applications) due to the need for an expensive setup phase that involves four communication rounds to establish shared randomness and pairwise keys in every FL training round (Ma et al., 2023).
In (Wang et al., 2023), Wang et al. proposed a single-mask secure aggregation protocol based on the Decisional Composite Residuosity (DCR) assumption, incorporating non-interactive zero-knowledge (NIZK) proofs to enable result verification. However, as noted by Wu et al. in (Wu and Zhang, 2024), the protocol is vulnerable to integrity attacks, where local gradients or aggregation results, along with their corresponding authentication tags or proofs, can be tampered without being detected by the verifiers. In (Fazli Khojir et al., 2023), Khojir et al. introduced a secure aggregation protocol based on additive secret sharing. The protocol utilizes a three-layered architecture (i.e., clients, middle servers, and lead server), which reduces communication costs for users compared to other protocols that rely on masking processes using secret sharing, which requires distributing secret shares among users, such as those in (Xu et al., 2020; Guo et al., 2021; et al., 2021a; Liu et al., 2023; Bell et al., 2020; Eltaras et al., 2023). However, it is unable to handle user dropouts, which is essential for real-world applications. In (Guo et al., 2024), Guo et al. proposed MicroSecAgg, a secure aggregation protocol designed to improve efficiency in single-server federated learning. However, Zhang et al. (Zhang et al., 2024) later identified a privacy vulnerability in MicroSecAgg, demonstrating that predictable masking values could be exploited to compromise user privacy.
Recently, in (Ma et al., 2023) and (Behnia et al., 2024), Ma et al. and Behnia et al. proposed “Flamingo” and “e-SeaFL”, respectively– multi-round, single-server secure aggregation protocols that do not require an initialization setup phase for each training round (a single setup phase suffices). These protocols significantly reduce computation and communication overhead compared to repeated setup-based approaches such as (et al., 2017; Xu et al., 2020; Guo et al., 2021; et al., 2021a; Liu et al., 2023; Bell et al., 2020; et al., 2024; Fu et al., 2024). However, as discussed in our security analysis in Appendix A, both Flamingo and e-SeaFL lack support for both forward and backward secrecy, which are important for single-setup secure aggregation protocols. Furthermore, they are not designed to operate in dynamic environments where new users may join or drop the training session at different times. Both protocols require prior knowledge of all participating users before the initial setup phase, due to the need for pairwise key generation.
Table 1 and Table 2 present a comparative analysis of our protocol against prominent masking-based secure aggregation protocols. Earlier protocols such as SecAgg (et al., 2017), SecAgg+ (Bell et al., 2020), and FastSecAgg (Kadhe et al., 2020) support forward and backward secrecy but require multiple setup rounds and lack support for dynamic user participation and model verification [^2] Model verification refers to the process of ensuring the integrity, authenticity, and correctness of the aggregated global model.. More recent protocols, including Flamingo (Ma et al., 2023) and e-SeaFL (Behnia et al., 2024), reduce setup complexity and introduce model verification capabilities, yet they do not offer forward/backward secrecy or support for dynamic users. In contrast, our protocol is the only single-setup protocol that simultaneously provides forward and backward secrecy, dynamic user participation, and model verification, while maintaining low communication and computational overhead as shown in Table 2333In terms of asymptotic computation and communication overhead, our protocol is comparable to e-SeaFL (Behnia et al., 2024). However, in practice, it achieves better computational efficiency, as e-SeaFL relies on relatively expensive elliptic curve operations and requires more frequent PRF evaluations. In contrast, our protocol leverages lightweight modular arithmetic. Detailed empirical performance results are provided in Section 8..
3. Forward and Backward Secrecy in Single-Setup Secure Aggregation
Forward secrecy (FS) and backward secrecy (BS) are essential privacy guarantees in secure aggregation based FL systems, especially in long-running deployments. While single-setup secure aggregation protocols like e-SeaFL (Behnia et al., 2024) and Flamingo (Ma et al., 2023) offer efficiency benefits, they lack FS and BS due to their reliance on static long-term secrets. This limitation poses a significant privacy risk: if a long-term secret is compromised, an adversary can reconstruct all past and future model updates in that training session. We provide a detailed case study of this vulnerability in Appendix A, including the mask derivation process and practical implications.
4. Preliminaries
@@
In this section, we briefly introduce the core concepts of symmetric homomorphic encryption and digital signatures, which are integral to the design of our protocol.
4.1. Symmetric Homomorphic Encryption
@@
Our protocol employs a key negation method to mask the user’s local model parameters. This method is based on the symmetric homomorphic encryption mechanism proposed in (Castelluccia et al., 2009). In this section, we will briefly describe the work of (Castelluccia et al., 2009). The key negation method will be explained in Section 6.1.
@@
$=$ $m_{i}+m_{j}+[(\cancel{\mathtt{k_{i}}+\mathtt{k_{j}}})-(\cancel{\mathtt{k_{i}}+\mathtt{k_{j}}})]\mod{q}$
(3) $=$ $m_{i}+m_{j}.$
The protocol described above has additive homomorphic properties (Castelluccia et al., 2009). The addition of two ciphertexts is illustrated in Equation 4.1. This property can also be extended to aggregate a set of ciphertexts using the corresponding aggregated secret keys, as shown in Equation 4.
(4) $\sum_{1\leq i\leq n}m_{i}=$ $\mathsf{Dec}_{\sum_{1\leq i\leq n}k_{i}}\left(\sum_{1\leq i\leq n}c_{i}\right)$
(5) $=$ $\sum_{1\leq i\leq n}c_{i}-\sum_{1\leq i\leq n}k_{i}\mod{q}.$
The described symmetric homomorphic encryption scheme is semantically secure (IND-CPA) if the secret keys $\mathtt{k_{i}}$, where $i\in\{1,2,\cdots,n\}$, are generated randomly and no keys are reused. Please refer to (Castelluccia et al., 2009) for detailed security proof.
4.2. Signature Scheme
Our protocol employs digital signatures to ensure the integrity and authenticity of messages exchanged between entities, leveraging a UF-CMA secure scheme as described in (et al., 2017). In general, a digital signature scheme consists of the following probabilistic polynomial-time (PPT) algorithms:
@@ 4.2. Signature Scheme
Figure 1. System Architecture
5. System Architecture, Threat Model, and Assumptions of Our Proposed Protocol
In this section, we present the system architecture, threat model, and security assumptions of our proposed protocol. We first start with the system architecture.
5.1. System Architecture
Figure 1 illustrates the system architecture of our protocol, which comprises three primary entities: users, intermediate servers, and the central server which we termed as Aggregator.
Users
They are the edge devices, such as smartphones and IoT devices, for cross-device FL [^4] Cross-device FL refers to a distributed machine learning method where many devices, such as smartphones or IoT devices, train a shared global model without sharing their local data., or organizations like hospitals, banks, universities, and government agencies for cross-silo FL [^5] Cross-silo FL involves a small number of organizations, such as hospitals or businesses, collaboratively training a model without sharing their private data. Unlike cross-device FL, cross-silo FL usually deals with higher-quality data from trusted institutions and is designed for scenarios with stricter privacy and security requirements.. These users generate and own data locally. They perform local model training on their own data, downloading the current global model from the aggregator, training it, and then sending the updated model parameters to the intermediate servers and aggregator. Before transmission, the parameters are masked, ensuring that only the aggregator can recover the final aggregated model parameter, thereby maintaining the confidentiality of individual users’ local model data.
Intermediate servers
They act as aggregation points between users and the main aggregator. They are to reduce communication overhead by combining model updates from multiple users before sending them to the aggregator. These servers receive model updates from users, aggregate the updates, and then forward the aggregated results to the aggregator for final processing. Additionally, they can help users detect model inconsistency attacks originating from the aggregator. In practical applications, fog nodes, edge servers, and third-party cloud services can function as intermediate servers, especially in edge computing or distributed system scenarios. By facilitating localized processing, these components reduce latency and bandwidth usage.
Aggregator
It is the central server that handles the entire FL process. It maintains the global model, coordinates the FL training rounds, and aggregates model updates directly from the intermediate servers and users. The aggregator initiates the training process by distributing the initial global model to users. After receiving aggregated updates from the intermediate servers, the aggregator combines these updates to improve the global model. It then sends the updated global model back to users for the next round of training.
5.2. Threat Model
@@
In the malicious model, we assume a more challenging scenario where both the aggregator and the intermediate servers may act maliciously, attempting to compromise the individual users’ local training models. In this scenario, we assume that the malicious aggregator and the malicious intermediate servers can collude with up to m2m-2 users, where $m$ is the total number of participating users in an FL round. Similarly, we assume that the malicious aggregator can collude with up to $n^{\prime}-1$ malicious intermediate servers in an FL round, where $n^{\prime}$ is the total number of participating intermediate servers. Additionally, our threat model considers model inconsistency attacks by the malicious aggregator, where the malicious aggregator provides different parameters to different users to exploit behavioral differences in the model updates, thereby inferring information on users’ datasets (Pasquini et al., 2022). We aim to detect such behavior by the malicious aggregator.
@@
We also consider the possibility of malicious users in our threat model. These users can collude with other users as well as with the aggregator and the intermediate servers to gain knowledge of other users’ local training models. However, we assume that up to m2m-2 users can collude, meaning at least two users must remain honest at all times. Without this assumption, the users could collude with the aggregator to obtain the local training model update of a targeted user.
5.3. Assumption
We make some assumptions in designing our protocol. First, we assume that the aggregator and intermediate servers remain online throughout the training process. Second, each entity possesses a public-private key pair managed via a Public Key Infrastructure (PKI). Third, all users are uniquely identifiable. Fourth, authenticated and private communication channels exist between users and intermediate servers, users and the aggregator, and between intermediate servers and the aggregator.
Figure 2. Sample Key Negation Mechanism